NatalieBM:

Sam and Bruno were computing how many kilometers they rode in the 3 bike trips they took last month. In order, they rode 45.7, 40.9, and 38 miles. Their solutions are shown. Sam’s Solution Bruno’s Solution 45.7 + 40.9 + 38.0 = 124.6 miles. 124.6 times 0.621 = 77.38 kilometers. 45.7 + 40.9 + 38.0 = 124.6 miles. 124.6 divided by 0.621 = 200.64 kilometers. If one kilometer equals 0.621 miles, which is the best solution to find the total number of kilometers they rode? Why is it correct? Bruno is correct. When going from a larger unit to a smaller unit you need to divide. Bruno is correct. When going from a smaller unit to a larger unit you need to divide. Sam is correct. When going from a larger unit to a smaller unit you need to multiply. Sam is correct. When going from a smaller unit to a larger unit you need to multiply.

smackz:

its b cause sam does 124.6 × 0.621 = 77.38 kilometers. 124.6 × 0.621 = 77.38 kilometers. bruno did 45.7 + 40.9 + 38.0 = 124.6 miles. 124.6 divided by 0.621 = 200.64 kilometers. now if u remember 1 kilometer = 0.621 miles. so therefore bruno is correct When going from a larger unit to a smaller unit you need to divide

NatalieBM:

ok

smackz:

i had to re do it vocal said i plagarized